Directions: Solve
25(V+1)^2-49=0

Answer by mananth(16949)

25(V+1)^2-49=0
25(V+1)^2=49
(V+1)^2=49/25
taking sqrt of the equation
v+1=7/5
5v+5=7
5v=7-5
5v=2
v=2/5